OpenStudy (anonymous):

The length of a rectangle is 2 feet less than three times the width. If the area of the rectangle is 40 square feet, find the dimensions.

OpenStudy (anonymous):

A=L*W A-Area L-Length W-Width

OpenStudy (anonymous):

L+2=3W

OpenStudy (anonymous):

Doesn't the 40 have something to do with the equation?

OpenStudy (anonymous):

A=(3W-2)*W=40 3W^2-2W=40

OpenStudy (anonymous):

3W^2-2W-40=0 w=4

OpenStudy (anonymous):

so L=10

OpenStudy (anonymous):

Can you elaborate more on how you got w

OpenStudy (anonymous):

\[ax ^{2}+bx+c=0\]

OpenStudy (anonymous):

\[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

substitute to the formula, here we have W instead of x

OpenStudy (anonymous):

Is the first one you showed completing the square?

OpenStudy (anonymous):

nope, it's the condition for the second one formula

OpenStudy (anonymous):

whenever you have an equation of such kind, you will apply that formula to find x

OpenStudy (anonymous):

Oh okay got it. Your just putting it in descending order to do the quadratic formula

OpenStudy (anonymous):

i disregarded negative value for width, as width can not be negative

OpenStudy (anonymous):

Okay thanks